OpenStudy (anonymous):

solve this equation two ways: 5x+20=5

OpenStudy (anonymous):

1. 5x + 20 = 5 5x = -15 x = -3

OpenStudy (anonymous):

That is one way

syed98 (syedmohammed98):



OpenStudy (anonymous):

The other way is 2. 5x + 20 = 5 x + 20 = 1 x = -19

OpenStudy (anonymous):

@Daina900 is what I wrote what you are looking for or are you looking for something different.

OpenStudy (idealist10):

5x+20=5 5x=5-20 5x=-15 x=-15/5 x=-3 Answer: -3